OpenStudy (cenaida):

Solve the following system of equations. x + y = 14 -x + y = 16 A. (1, 15) B. (1, -15) C. (-1, 15) D. (6, 6)

OpenStudy (luigi0210):

Add the two equations, that will cancel out the x's and leave y

OpenStudy (cenaida):

I don't understand...

OpenStudy (anonymous):

Solve for one of the variable in either equation. Then plug in the variables answer into the other equation and you will solve for a variable and find the other variable as well and you will come up with a point. Want me to evaluate?

OpenStudy (luigi0210):

Oh Cenaida :/

OpenStudy (luigi0210):

You're not allowed to do their work for them, Gerardo.

jimthompson5910 (jim_thompson5910):

if you're still stuck, then I'll get you started Add the equations to get x + y = 14 -x + y = 16 ------------ 0x+2y = 30 so 2y = 30 which means y = ???

OpenStudy (cenaida):

I don't know :/ 1?

jimthompson5910 (jim_thompson5910):

2 times y = 30 so y must equal what number?

OpenStudy (cenaida):

15~ :D

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now that you know that y = 15, use this to find x

OpenStudy (cenaida):

and how do I find x?

jimthompson5910 (jim_thompson5910):

plug in y = 15 into any equation, then solve for x so let's pick on the first equation x + y = 14 x + 15 = 14 ... replace y with 15 (since y = 15) now solve for x

OpenStudy (cenaida):

i don't know how to solve for x......

jimthompson5910 (jim_thompson5910):

x is some number

jimthompson5910 (jim_thompson5910):

so saying x+15 = 14 is the same as saying "some number" + 15 = 14 what replaces "some number"?

jimthompson5910 (jim_thompson5910):

to find out, undo that +15 and subtract 15 from both sides

OpenStudy (cenaida):

-1?

jimthompson5910 (jim_thompson5910):

good, x = -1

jimthompson5910 (jim_thompson5910):

so we found that x = -1 and y = 15 which means that the ordered pair solution is (-1,15)

OpenStudy (cenaida):

ok thanks jim!

jimthompson5910 (jim_thompson5910):

yw